OpenStudy (starryknight):

which of the following is not a proportion 2/6=7/21 5/6=20/24 8/0.4=6/0.3 3/4=4/3

OpenStudy (anonymous):

Well, a proportion is when you have two fractions, and when you cross multiply the same number comes up. For ex. 2/4 = 4/ 8. 2 x 8 = 16 and 4 x 4 = 16 This IS a proportion. Now with your choices, just cross multiply and see if they are equal to find your answer!

OpenStudy (starryknight):

thanx i was confused ^^

OpenStudy (anonymous):

You're welcome! Do you know now how to find the answer now, or do you need further help with that?

OpenStudy (starryknight):

Im good now